OpenStudy (anonymous):

Complex Derivatives !! if y = ((6x-3)^4)/((3x+4)^(1/3)) what is dy/dx ?

OpenStudy (anonymous):

\[y = \frac{ (6x-3)^4 }{ \sqrt[3]{3x+4} }\]

OpenStudy (anonymous):

okay so let me show you what i got and then tell me what iv done wrong

OpenStudy (danjs):

did you move the root to the top as -1/3 power, and apply the product rule

OpenStudy (anonymous):

!!! no i was doing the quotient rule like the fool i am!!! very clever!!

OpenStudy (anonymous):

okay let me start over

OpenStudy (danjs):

either is good, it isnt too much work

OpenStudy (anonymous):

okay so this is what iv got so far \[(24(6x-3)^3)((3x+4)^{-1/3}) + ((6x-3)^4)(-(3x+4)^{-4/3})\]

OpenStudy (anonymous):

quick question : i cant distribute the 24 into the (6x-3) because its cubed right?

OpenStudy (danjs):

no

OpenStudy (anonymous):

so i can distribute the 24?

OpenStudy (danjs):

test and see if you forget algebra things... 24(6x - 3)^3 = the distributed value?

OpenStudy (danjs):

use any number for x

OpenStudy (anonymous):

okay sorry lost internet connection,

OpenStudy (danjs):

maybe do the quotient rule and see if that is easier

OpenStudy (anonymous):

not sure where to go now, do i just bring down the negative exponents?

OpenStudy (anonymous):

\[(24(6x-3)^3 + (6x-3)^4)/(-(3x+4)^{4/3}+(3x+4)^{1/3})\]

OpenStudy (anonymous):

??

OpenStudy (danjs):

that is the derivative, done..

OpenStudy (danjs):

i am sure they want it all simplified though i assume

OpenStudy (anonymous):

yep it seems so , it dosent appear to be any of the answer choices

OpenStudy (danjs):

you can move the - exponents , but you can not do that, a * b^-1 + c*d^-1 is not same as (a +c) / (b + d)

OpenStudy (danjs):

rather it is just (a/b) + (c/d)

OpenStudy (anonymous):

okay don't worry about the simplification it is just tedious work i think iv got it , thank you very much for your help

OpenStudy (danjs):

notice the parenthesis value appears in both terms, for both the different parenthesis, maybe try some factoring

OpenStudy (danjs):

\[(24(6x-3)^3)((3x+4)^{-1/3}) + ((6x-3)^4)(-(3x+4)^{-4/3})\]

OpenStudy (danjs):

or, put the negative exponents back to the bottom, then you would have to do the common denominator to add the two fractions

OpenStudy (danjs):

\[\frac{ 24(6x-3)^3 }{ (3x+4)^{1/3} } + \frac{ (6x-3)^4 }{ (3x+4)^{4/3}}\]

OpenStudy (danjs):

need to multiply first term by (3x+4)^{3/3} / (3x+4)^{3/3}

OpenStudy (anonymous):

you mean 4/3 for your exponent right?

OpenStudy (anonymous):

not that it matters much now but we did it completely wrong, we were suppose to multiply both sides by Ln at the beginning and work from there

OpenStudy (danjs):

yes, i just skipped a couple things, the net effect is what i said the common denominator to multiply everything by is the product of the two, so you get (3x+4)^{5/3} that puts the denominator to 5/3 power, and adds a factor of (3x+4)^{1/3} that can factor from the top two terms, so overall you are just multiplying the first fraction by (3x+4)^(3/3)

OpenStudy (danjs):

\[\frac{ (3x+4)*24*(6x-3)^2 + (6x-3)^4 }{ (3x+4)^{4/3} }\]